OpenStudy (help_needed_cudi):

PLEASE HELP!! I give medals and i need each step on how to get there and i have more problems i need help on. 3/√2−3

OpenStudy (anonymous):

I'm pretty sure its already simplified

OpenStudy (help_needed_cudi):

thanks i thought so also but my teacher put it on my test do you think you could help me with more?

OpenStudy (help_needed_cudi):

6. 􀶥𝑘55√18𝑘 is the next one and i need help to check 3 questions....@severus4ever

OpenStudy (help_needed_cudi):

imma re due that

OpenStudy (anonymous):

\[\frac{ 3 }{\sqrt{2} -3 }\times \frac{ \sqrt{2} +3}{ \sqrt{2} +3}=\frac{ 3\left( \sqrt{2} +3\right) }{ 2-9 }\] \[=\frac{ 3\sqrt{2} +9}{ -7 }\]

OpenStudy (help_needed_cudi):

thanks my internet had went out but thank you!!!!

OpenStudy (anonymous):

yw